# Classification of nilpotent almost abelian Lie groups admitting left-invariant complex or symplectic structures

### v1

@@ Abstract.
[…]
@@
Nilpotent almost abelian Lie group, complex structure, symplectic structure
@@ 1. Introduction
Nilmanifolds admitting geometric structures play an important role in differential geometry. They have provided examples and counterexamples to open questions and conjectures and have been extensively studied by several authors (see [1, 5, 6, 7, 13, 16, 31, 33, 35, 40], among many others). Most of the first known examples of compact symplectic manifolds which do not admit any Kähler structure are nilmanifolds (see, for instance, [15, 39]). It is well known that the only nilmanifolds admitting Kähler structures are tori [8, Theorem A].
A nilmanifold is a manifold of the form $\Gamma\backslash G$, where $G$ is a simply-connected nilpotent Lie group and $\Gamma\subset G$ is a lattice, that is, a co-compact discrete subgroup of $G$. In general, it is not easy to determine whether a given Lie group $G$ admits a lattice. In the nilpotent case there is a well-known criterion due to Malcev [32], namely, a simply connected nilpotent Lie group has a lattice if and only if its Lie algebra admits a basis with respect to which the structure constants are rational. An important feature of nilmanifolds is that their de Rham cohomology can be computed in terms of left-invariant forms. Indeed, Nomizu [34] proved that there is an isomorphism $H^{*}_{dR}(\Gamma\backslash G)\cong H^{*}(\mathfrak{g})$. In particular, $H^{*}_{dR}(\Gamma\backslash G)$ does not depend on the lattice $\Gamma$.
In [19] it was shown that bi-invariant complex structures are closely related to Hermitian manifolds which are Chern-flat and quasi-Kähler. In fact, every compact manifold satisfying these conditions is isometric to the quotient of a complex 2-step nilpotent Lie group $G$ by a co-compact lattice $\Gamma$. Complex structures on 2-step nilpotent Lie groups are also related to other geometric structures on such groups. Indeed, it was shown in [3] that a 2-step nilpotent Lie group $G$ admits a left-invariant Killing-Yano 2-form if and only $G$ admits a bi-invariant complex structure.
The study of left-invariant complex or symplectic structures on nilpotent Lie groups is an active field of research and several results are known in low dimensions (see for instance, [8, 9, 10, 11, 12, 15, 17, 18, 20, 21, 22, 24, 27, 28, 36, 37, 38]). The classification of nilpotent Lie groups admitting left-invariant complex or symplectic structures is an interesting problem. Although there are many contributions to the subject, it is considered a wild problem and in general it is far from being solved.
In this article we focus on a family of nilpotent Lie groups, namely, the nilpotent almost abelian ones. A Lie group $G$ is called almost abelian if its Lie algebra $\mathfrak{g}$ has a codimension-one abelian ideal $\mathfrak{h}$, that is, given $e_{0}\notin\mathfrak{h},$ $\mathfrak{g}$ decomposes as $\mathfrak{g}={\mathbb{R}}e_{0}\ltimes\mathfrak{h}$. Such a Lie algebra will be also called almost abelian. Fix a basis $\mathcal{B}$ of $\mathfrak{h}$ and let $A$ be the matrix of $\operatorname{ad}_{e_{0}}|_{\mathfrak{h}}$ with respect to $\mathcal{B}$. If $\dim\mathfrak{h}=d$, then $\mathfrak{g}$ is isomorphic to $\mathfrak{g}_{A}:={\mathbb{R}}e_{0}\ltimes_{A}{\mathbb{R}}^{d}$, with $[e_{0},v]=Av$, where we consider $v\in{\mathbb{R}}^{d}$ as a column vector. Accordingly, the corresponding simply connected Lie group $G_{A}$ is a semidirect product $G_{A}={\mathbb{R}}\ltimes_{\phi}{\mathbb{R}}^{d}$, where ${\mathbb{R}}$ acts on ${\mathbb{R}}^{d}$ by $\phi(t)=e^{tA}$, that is, group multiplication in $G_{A}$ is given by:
@@
The almost abelian Lie algebra $\mathfrak{g}_{A}$ is nilpotent if and only if $A$ is nilpotent. Isomorphism classes of $2n$-dimensional nilpotent almost abelian Lie algebras are parametrized by conjugacy classes of $(2n-1)\times(2n-1)$ nilpotent matrices. These are in one-to-one correspondence with partitions of $12n-$ 1, due to the Jordan decomposition theorem. In our classification results, Theorems 3.10 and 4.10, we determine which partitions of $12n-$ 1 correspond to $2n$-dimensional nilpotent almost abelian Lie algebras admitting complex or symplectic structures. The proof of Theorem 3.10 essentially relies on the computation of the Jordan normal form of a nilpotent matrix in $M_{J_{m}}(2m,{\mathbb{R}})$ (see (6)). On the other hand, the theory of nilpotent orbits in $\mathfrak{sp}(2m,{\mathbb{R}})$ [14] (see also [26, §1.4 and §1.6]) plays a crucial role in the proof of Theorem 4.10. As a consequence of Theorems 3.10 and 4.10 we obtain Theorem 4.11, which asserts that if $\mathfrak{g}_{A}$ admits a complex structure, then it necessarily admits a symplectic structure. This does not hold for arbitrary nilpotent Lie algebras (see Example 4.12).
Let $\mathfrak{g}_{A}$ be a nilpotent almost abelian Lie algebra and $G_{A}$ the corresponding simply connected Lie group. It follows from Malcev’s criterion that $G_{A}$ always has lattices. Given a lattice $\Gamma\subset G_{A}$, any left-invariant symplectic structure on $G_{A}$ induces a symplectic structure on $\Gamma\backslash G_{A}$. Similarly, any left-invariant complex structure on $G_{A}$ induces a complex structure on $\Gamma\backslash G_{A}$. As a consequence of our main theorems we obtain infinitely many nilmanifolds equipped with complex or symplectic structures, possibly both. According to [8, Theorem A], none of these nilmanifolds admits a Kähler structure.
Next, we summarize the remaining results obtained in this article. Proposition 3.6 shows that every even dimensional 2-step nilpotent almost abelian Lie algebra admits a complex structure. However, this is no longer true for the 3-step nilpotent almost abelian Lie algebras. In Proposition 3.8 we study this case. On the other hand, Theorem 4.6 shows that any even dimensional $k$-step nilpotent almost abelian Lie algebra admits a symplectic structure, for $k=2,3,4$.
As an application of our main results, we classify $8$-dimensional nilpotent almost abelian Lie algebras admitting complex or symplectic structures (see Corollaries 5.1 and 5.2, respectively). In Corollaries 5.3 and 5.4, we analyze the existence of complex or symplectic structures in dimension $10$, where we find the only nilpotent almost abelian Lie algebra which admits neither complex nor symplectic structures in this dimension (see Example 4.7). Moreover, this is the lowest dimension and the lowest step of nilpotency where this occurs.
In Corollaries 5.6 and 5.7 we study $k$-step nilpotent almost abelian Lie algebras for $k\geq 5$. We find upper bounds for the dimension of the Lie algebra to ensure the existence of a symplectic structure.
Finally, in Corollary 5.8, we exhibit families of nilpotent almost abelian Lie algebras admitting both, complex and symplectic structures, or symplectic but no complex structure, or neither complex nor symplectic structure.
Acknowledgements. The authors are grateful to Leandro Cagliero for helpful comments.
@@ 2. Preliminaries
Given a Hermitian structure $(j,{\langle\cdot,\cdot\rangle})$ on $\mathfrak{g}$, the Kähler form associated to $(j,{\langle\cdot,\cdot\rangle})$ is defined by:
$\omega_{j}(x,y)=\langle jx,y\rangle,\qquad\quad x,y\in\mathfrak{g}.$
The Hermitian structure $(j,{\langle\cdot,\cdot\rangle})$ is said to be Kähler when $\omega_{j}$ is closed.
If $G$ is a connected Lie group with Lie algebra $\mathfrak{g}$, any complex structure on $\mathfrak{g}$ induces a left-invariant complex structure on $G$. Similarly, a symplectic (resp. Kähler) structure on $\mathfrak{g}$ induces a left-invariant symplectic (resp. Kähler) structure on $G$.
@@
In view of the previous remark, we state and prove the following lemma.
@@
$\text{rk}\,(A^{n})=\begin{cases}\begin{array}[]{cc}\text{rk}\,(B^{n}),&\text{if}\hskip 7.11317ptB^{n-1}v\in{\rm Im}\,B^{n}\\ \text{rk}\,(B^{n})+1,&\text{if}\hskip 7.11317ptB^{n-1}v\notin{\rm Im}\,B^{n}.\end{array}\end{cases}$
@@
$\text{rk}\,(C^{n})=\begin{cases}\begin{array}[]{cc}\text{rk}\,(B^{n}),&\text{if}\hskip 7.11317ptB^{n-1}v\in{\rm Im}\,B^{n}\\ \text{rk}\,(B^{n})+1,&\text{if}\hskip 7.11317ptB^{n-1}v\notin{\rm Im}\,B^{n}.\end{array}\end{cases}$
@@
Throughout this article, we use the following notation. The vector $\epsilon_{\ell}$ denotes the $\ell$th vector of the canonical basis of ${\mathbb{R}}^{d}$, for $\ell=1,\dots,d$. Given two matrices $M_{i}\in M(d_{i},\mathbb{R})$, for $i=1,2$, we denote by $M_{1}\oplus M_{2}$ the matrix $\begin{pmatrix}M_{1}&0\\ 0&M_{2}\end{pmatrix}\in M(d_{1}+d_{2},\mathbb{R})$ and we use a similar notation for 3 or more matrices. In particular, given a matrix $M$, we write $M^{\oplus p}$ for the matrix $M\oplus\cdots\oplus M$ ($p$ copies). On the other hand, $\mathcal{J}_{n}$ denotes the nilpotent Jordan block of size $n$:
(1) $\mathcal{J}_{n}=\begin{pmatrix}0&0&\cdots&\cdots&0\\ 1&0&&&0\\ 0&\ddots&\ddots&&\vdots\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ 0&\cdots&0&1&0\end{pmatrix}.$
Finally, $0_{s}$ denotes the null $s\times s$ matrix.
The idea of the proof of the next result is contained in [4, Theorem 3.4].
Proposition 2.4.
Let $v\in\mathbb{R}^{m}$ and let $C=\begin{pmatrix}0&0\\ v&B\end{pmatrix}\in M(m+1,{\mathbb{R}})$, where $B=\mathcal{J}_{m_{1}}\oplus\cdots\oplus\mathcal{J}_{m_{r}}\oplus 0_{s}$, with $m_{1}\geq\cdots\geq m_{r}\geq 2$ and $s\geq 0$. Then $C$ is conjugate to one of the following matrices:
  1. ${\rm(a)}$
    $0_{1}\oplus B$;
  2. ${\rm(b)}$
    $\mathcal{J}_{m_{1}}\oplus\cdots\oplus\mathcal{J}_{m_{d-1}}\oplus\mathcal{J}_{m_{d}+1}\oplus\mathcal{J}_{m_{d+1}}\oplus\cdots\oplus\mathcal{J}_{m_{r}}\oplus 0_{s}$, for $d\in\{1,\dots,r\}$;
  3. ${\rm(c)}$
Proof.
Assume first that $v\in{\rm Im}\,B$. Hence, by Lemma 2.3, $C$ is conjugate to $0_{1}\oplus B$.
Now, let $t=m_{1}+\cdots+m_{r}$ and we rename the elements of the canonical basis of $\mathbb{R}^{m}$ as $\{\epsilon^{1}_{1},\dots,\epsilon^{1}_{m_{1}},\dots,\epsilon_{1}^{r},\dots,\epsilon_{m_{r}}^{r},\epsilon_{t+1},\dots,\epsilon_{t+s}\}$. We can write $v$ as
@@ 3. ${\rm(c)}$
where $v_{i}\in\mathbb{R}$, $w\in{\rm Im}\,B$ and $w_{0}\in\text{span}\{\epsilon_{t+1},\ldots,\epsilon_{t+s}\}\subseteq\operatorname{Ker}\,B$. By Lemma 2.3, we may assume that $w=0$.
If $v\notin{\rm Im}\,B$, then either there exists $i$ such that $v_{i}\neq 0$ or $v=w_{0}\neq 0$.
Assume first that there exists $i$ such that $v_{i}\neq 0$ and set $d=\text{min}\{i:v_{i}\neq 0\}$. Let
@@
Since $C$ and $A$ are both nilpotent matrices such that $C^{m_{1}+1}=A^{m_{1}+1}=0$, it suffices to show that $\dim(\operatorname{Ker}\,C^{n})=\dim(\operatorname{Ker}\,A^{n})$, for $1\leq n\leq m_{1}$.
In general, we observe that
(2) $\dim(\operatorname{Ker}\,\mathcal{J}^{n}_{m_{i}})=\begin{cases}\begin{array}[]{cc}m_{i},&\text{if }m_{i}<n\\ n,&\text{if }m_{i}\geq n\end{array}\end{cases}$
Since $v\notin{\rm Im}\,B$,

### v2

# Classification of almost abelian Lie groups admitting left-invariant complex or symplectic structures

@@ Abstract.
We classify the almost abelian Lie algebras $\mathfrak{g}_{A}={\mathbb{R}}e_{0}\ltimes_{A}{\mathbb{R}}^{2n-1}$ admitting complex or symplectic structures. The matrix $A\in M(2n-1,{\mathbb{R}})$ encodes the adjoint action of $e_{0}$ on the abelian ideal ${\mathbb{R}}^{2n-1}$, and the existence of complex or symplectic structures on $\mathfrak{g}_{A}$ imposes restrictions on the Jordan normal form of $A$. The classification essentially reduces to the case when $A$ is nilpotent, so we start by considering this case. It turns out that if $A$ is nilpotent and $\mathfrak{g}_{A}$ admits a complex structure, then $\mathfrak{g}_{A}$ necessarily admits a symplectic structure. This is not true in general when $A$ is non-nilpotent. Finally, several consequences of the classification theorems are obtained.
@@
Almost abelian Lie algebra, complex structure, symplectic structure
@@ 1. Introduction
Nilmanifolds, and more generally, solvmanifolds admitting geometric structures play an important role in differential geometry. They have provided examples and counterexamples to open questions and conjectures and have been extensively studied by several authors (see [1, 2, 4, 5, 6, 7, 12, 10, 11, 14, 17, 22, 28, 34, 36, 37, 44], among many others). Most of the first known examples of compact symplectic manifolds which do not admit any Kähler structure are nilmanifolds (see, for instance, [16, 43]). It is well known that the only nilmanifolds admitting Kähler structures are tori [8, Theorem A].
The study of left-invariant complex or symplectic structures on solvable Lie groups is an active field of research and several results are known in low dimensions (see for instance, [8, 9, 10, 11, 12, 13, 16, 18, 19, 20, 21, 22, 26, 25, 24, 30, 31, 40, 38, 39, 41, 42]). The classification of solvable Lie groups admitting left-invariant complex or symplectic structures is an interesting problem (see for instance [29]). Although there are many contributions to the subject, it is considered a wild problem and it is far from being solved in general.
In this article we focus on a family of solvable Lie groups, namely, the almost abelian ones. A Lie group $G$ is called almost abelian if its Lie algebra $\mathfrak{g}$ has a codimension-one abelian ideal $\mathfrak{h}$, that is, given $e_{0}\notin\mathfrak{h},$ $\mathfrak{g}$ decomposes as $\mathfrak{g}={\mathbb{R}}e_{0}\ltimes\mathfrak{h}$. Such a Lie algebra will be also called almost abelian. Fix a basis $\mathcal{B}$ of $\mathfrak{h}$ and let $A$ be the matrix of $\operatorname{ad}_{e_{0}}|_{\mathfrak{h}}$ with respect to $\mathcal{B}$. If $\dim\mathfrak{h}=d$, then $\mathfrak{g}$ is isomorphic to $\mathfrak{g}_{A}:={\mathbb{R}}e_{0}\ltimes_{A}{\mathbb{R}}^{d}$, with $[e_{0},v]=Av$, where we consider $v\in{\mathbb{R}}^{d}$ as a column vector. Accordingly, the corresponding simply connected Lie group $G_{A}$ is a semidirect product $G_{A}={\mathbb{R}}\ltimes_{\phi}{\mathbb{R}}^{d}$, where ${\mathbb{R}}$ acts on ${\mathbb{R}}^{d}$ by $\phi(t)=e^{tA}$, that is, group multiplication in $G_{A}$ is given by:
@@
In order to prove the classification theorems we need to obtain the conjugacy classes of certain $(2n-1)\times(2n-1)$ matrices which characterize the almost abelian Lie algebras admitting complex or symplectic structures.
We start by considering the nilpotent case (see Sections §4.2 and §5.2). Isomorphism classes of $2n$-dimensional nilpotent almost abelian Lie algebras are parametrized by conjugacy classes of $(2n-1)\times(2n-1)$ nilpotent matrices. These are in one-to-one correspondence with partitions of $12n-$ 1, due to the Jordan normal form of nilpotent matrices. In our classification results, Theorems 4.10 and 5.10, we determine which partitions of $12n-$ 1 correspond to $2n$-dimensional nilpotent almost abelian Lie algebras admitting complex or symplectic structures. As a consequence of Theorem 4.10 and Theorem 5.10 we obtain Theorem 5.13, which asserts that if $\mathfrak{g}_{A}$ admits a complex structure, then it necessarily admits a symplectic structure. This does not hold for arbitrary nilpotent Lie algebras (see Example 5.17).
The general case follows by applying suitably the results of the nilpotent case. The classification is carried out by making use of the description of conjugacy classes of matrices in $M_{J_{m}}(2m,{\mathbb{R}})$ or $\mathfrak{sp}(2m,{\mathbb{R}})$ (see Proposition 4.5 and Theorem 5.6, respectively). The main results are Theorem 4.17 and Theorem5.26, and from these we obtain the classification in the complex and symplectic case, respectively.
Several consequences of our main results are obtained throughout the paper.
Acknowledgements. The authors are grateful to Leandro Cagliero for insightful discussions on nilpotent orbits in the symplectic Lie algebra, and to Jorge Lauret for suggesting to explore the validity of Theorem 6.4. We owe special thanks to Marco Freibert for drawing our attention to reference [6] and for encouraging us to extend our results to the solvable case, which led to significant improvements in the article.
@@ 2. Preliminaries
If $G$ is a connected Lie group with Lie algebra $\mathfrak{g}$, any complex structure on $\mathfrak{g}$ induces a left-invariant complex structure on $G$. In this way, $G$ becomes a complex manifold such that left translations are holomorphic diffeomorphisms. Similarly, a symplectic structure on $\mathfrak{g}$ induces a left-invariant symplectic structure on $G$.
@@
3. Useful linear algebra results
Given two matrices $M_{i}\in M(d_{i},\mathbb{R})$, for $i=1,2$, we denote by $M_{1}\oplus M_{2}$ the matrix $\begin{pmatrix}M_{1}&0\\ 0&M_{2}\end{pmatrix}\in M(d_{1}+d_{2},\mathbb{R})$ and we use a similar notation for 3 or more matrices. In particular, given a matrix $M$, we write $M^{\oplus p}$ for the matrix $M\oplus\cdots\oplus M$ ($p$ copies). Finally, $0_{s}$ denotes the null $s\times s$ matrix and $I_{t}$ denotes the identity matrix of size $t\times t$.
Let $C\in M(r,{\mathbb{R}})$, we denote by $\text{spec}\,(C)$ the set of distinct eigenvalues of $C$ and by $\text{spec}_{\mathbb{R}}(C)$ the set of distinct real eigenvalues of $C$. For $a\in{\mathbb{R}}$ and $\alpha\in{\mathbb{C}}$ with Im$(\alpha)>0$, we consider the following sets of matrices:
$M_{a}(r,{\mathbb{R}})$ ::=:= $\{C\in M(r,{\mathbb{R}}):\,\text{spec}\,(C)=\{a\}\},$
$M_{\alpha}(r,{\mathbb{R}})$ ::=:= $\{C\in M(r,{\mathbb{R}}):\,\text{spec}\,(C)=\{\alpha,\overline{\alpha}\}\}.$
Note that when $\alpha\in{\mathbb{C}}$ with Im$(\alpha)>0$, $r$ is necessarily even.
Given $\lambda\in{\mathbb{C}}$ and $n\geq 2$, $\mathcal{J}_{n}(\lambda)$ denotes the Jordan block of size $n\times n$ :
$\mathcal{J}_{n}(\lambda)=\begin{pmatrix}\lambda&0&\cdots&\cdots&0\\ 1&\lambda&&&0\\ 0&\ddots&\ddots&&\vdots\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ 0&\cdots&0&1&\lambda\end{pmatrix}.$
When $\lambda=0$, we simply denote $\mathcal{J}_{n}$ instead of $\mathcal{J}_{n}(0)$.
For each $m\in{\mathbb{N}}$, we denote by $\bigwedge_{m}$ the following set of tuples of integers:
$\begin{split}{\bigwedge}_{m}:&=\{(n_{1},\ldots,n_{k};p_{1},\ldots,p_{k};t):\sum_{i}n_{i}p_{i}+t=m,\\ &\qquad k\geq 1,\;n_{1}>\dots>n_{k}\geq 2,\;p_{i}>0\;\forall i,\;t\geq 0\}\cup\{(m)\}.\end{split}$
Let $A\in M_{b}(m,{\mathbb{R}})$, $b\in{\mathbb{R}}$. Then $A$ is conjugate to either
(1) $\begin{cases}\mathcal{J}_{n_{1}}(b)^{\oplus p_{1}}\oplus\dots\oplus\mathcal{J}_{n_{k}}(b)^{\oplus p_{k}}\oplus bI_{t},\quad(n_{1},\ldots,n_{k};p_{1},\ldots,p_{k};t)\in{\bigwedge}_{m},\\ \text{or}\\ bI_{m}.\end{cases}$
We associate to $A$ the following tuple $\Lambda(A)$:
(2) $\Lambda(A)=(n_{1},\ldots,n_{k};p_{1},\ldots,p_{k};t),\quad\text{ or }\quad\Lambda(A)=(m),\text{ respectively.}$.
Given $b,d\in{\mathbb{R}}$ where $d>0$, let
(3) $\mathcal{C}(b,d)=\begin{pmatrix}b&d\\ -d&b\end{pmatrix}.$
For $n\geq 2$ we define
(4) $\mathcal{C}_{n}(b,d)=\begin{pmatrix}\mathcal{C}(b,d)&&&&\\ I_{2}&\mathcal{C}(b,d)&&&\\ &I_{2}&\ddots&&\\ &&\ddots&\ddots&\\ &&&I_{2}&\mathcal{C}(b,d)\end{pmatrix}\in M(2n,{\mathbb{R}}).$
Let $\alpha=b+id$ with $d>0$ and let $A\in M_{\alpha}(2m,{\mathbb{R}})$. Then $A$ is conjugate to either
(5) $\begin{cases}\mathcal{C}_{n_{1}}(b,d)^{\oplus p_{1}}\oplus\dots\oplus\mathcal{C}_{n_{k}}(b,d)^{\oplus p_{k}}\oplus\mathcal{C}(b,d)^{\oplus t},\quad(n_{1},\ldots,n_{k};p_{1},\ldots,p_{k};t)\in{\bigwedge}_{m},\\ \text{or}\\ \mathcal{C}(b,d)^{\oplus m}.\end{cases}$
We associate to $A$ the following tuple $\Lambda(A)$:
(6) $\Lambda(A)=(n_{1},\ldots,n_{k};p_{1},\ldots,p_{k};t),\quad\text{ or }\quad\Lambda(A)=(m),\text{ respectively.}$.
Let $C\in M(r,{\mathbb{R}})$ such that spec$\,(C)=\{\lambda_{1},\ldots,\lambda_{\ell},\lambda_{\ell+1},\overline{\lambda}_{\ell+1},\ldots,\lambda_{s},\overline{\lambda}_{s}\}$, where $\lambda_{j}\in{\mathbb{R}}$, for $j=1,\ldots,\ell$, and $\lambda_{j}\in{\mathbb{C}}$ with $\text{Im}(\lambda_{j})>0$, for $j=\ell+1,\ldots,s$.
Then $C$ is conjugate to
(7) $C(\lambda_{1})\oplus\cdots\oplus C(\lambda_{s}).$
where $C(\lambda_{j})\in M_{\lambda_{j}}(m_{j},{\mathbb{R}})$, $m_{1}+\cdots+m_{s}=r$ and
  • if $j=1,\ldots,\ell$, then $C(\lambda_{j})$ is one of the matrices given by (1) with $b=\lambda_{j}$ and $m=m_{j}$,
  • if $j=\ell+1,\ldots,s$, the matrix $C(\lambda_{j})$ is one of the matrices given by (5) with $\lambda_{j}=b_{j}+id_{j}$ and $2m=m_{j}$.
Through the paper, given a real square matrix $M$ and $\lambda\in\text{spec}_{\mathbb{R}}(M)$, we denote $m_{\lambda}$ the positive integer satisfying $M(\lambda)\in M(m_{\lambda},{\mathbb{R}})$.
The next lemma is a straightforward consequence of (7).
Lemma 3.1.
Given $C\in M(m,{\mathbb{R}})$ and $b\in\text{spec}_{\mathbb{R}}(C)$, then $C$ is conjugate to either $C(b)$ or $C(b)\oplus Q$, where $Q\in M(m-m_{b},{\mathbb{R}})$ and $b\notin\text{spec}\,(Q)$, depending on whether $m_{b}=m$ or $m_{b}<m$.
Remark 3.2.
Observe that if $b\in{\mathbb{R}}$ and $C\in M_{b}(m,{\mathbb{R}})$, then $\Lambda(C)=\Lambda(C-bI_{m})$, since $C(b)=(C-bI_{m})(0)+bI_{m}$.
Given $b\in{\mathbb{R}}$, we denote by $M_{b}(m,{\mathbb{R}})/GL(m,{\mathbb{R}})$ the set of conjugacy classes of matrices in $M_{b}(m,{\mathbb{R}})$. It follows from uniqueness of the Jordan normal form up to reordering that tuples corresponding to conjugate matrices in $M_{b}(m,{\mathbb{R}})$ are equal. Therefore, $\Lambda$ as in (2) induces a well defined map $\tilde{\Lambda}$:
(8) $\tilde{\Lambda}:M_{b}(m,{\mathbb{R}})/GL(m,{\mathbb{R}})\longrightarrow{\bigwedge}_{m},\qquad\quad[A]\mapsto\Lambda(A),$
where $[A]$ denotes the conjugacy class of $A\in M_{b}(m,{\mathbb{R}})$. The map $\tilde{\Lambda}$ is clearly a bijection between $M_{b}(m,{\mathbb{R}})/GL(m,{\mathbb{R}})$ and ${\bigwedge}_{m}$.
In view of Proposition 2.1, in order to study isomorphism classes of almost abelian Lie algebras admitting complex or symplectic structures, we need to determine the conjugacy classes of matrices of the form (12) and (17). We will consider separately the nilpotent and non-nilpotent case.
3.1. The nilpotent case
In view of Remark 2.2, we state and prove the following lemma.
Lemma 3.3.
Let $m\in{\mathbb{N}}$. The isomorphism classes of $(m+1)$-dimensional nilpotent almost abelian Lie algebras are parametrized by the set $\bigwedge_{m}$.
Proof.
According to Remark 2.2, the isomorphism classes of $(m+1)$-dimensional nilpotent almost abelian Lie algebras are in one-to-one correspondence with the set of conjugacy classes of nilpotent $m\times m$ matrices. This set is parametrized by $\bigwedge_{m}$ via the bijection $\tilde{\Lambda}$ defined in (8) for $b=0$, and the lemma follows. ∎
In the following results we study the Jordan normal form of nilpotent matrices of the form $\begin{pmatrix}0&0\\ v&B\end{pmatrix}\in M(m+1,{\mathbb{R}})$, where $v\in{\mathbb{R}}^{m}$ and $B\in M(m,{\mathbb{R}})$.
Proposition 3.4.
$B=\mathcal{J}_{m_{1}}\oplus\cdots\oplus\mathcal{J}_{m_{r}}\oplus 0_{s},\quad\text{ with }\;m_{1}\geq\cdots\geq m_{r}\geq 2,\;s\geq 0.$
Then $C$ is conjugate to one of the following matrices:
  1. ${\rm(a)}$
    $0_{1}\oplus B$,
  2. ${\rm(b)}$
    $\mathcal{J}_{m_{1}}\oplus\cdots\oplus\mathcal{J}_{m_{d-1}}\oplus\mathcal{J}_{m_{d}+1}\oplus\mathcal{J}_{m_{d+1}}\oplus\cdots\oplus\mathcal{J}_{m_{r}}\oplus 0_{s}$, for $d\in\{1,\dots,r\}$,
  3. ${\rm(c)}$
If $B=0_{m}$ and $v\neq 0$, then $C$ is conjugate to $\mathcal{J}_{2}\oplus 0_{m-1}$.
For the proof of the above proposition we need the next lemma.
@@ 3. ${\rm(c)}$
$\text{rk}\,(A^{n})=\begin{cases}\begin{array}[]{cc}\text{rk}\,(B^{n}),&\text{if}\hskip 7.11317ptB^{n-1}v\in{\rm Im}\,B^{n},\\ \text{rk}\,(B^{n})+1,&\text{if}\hskip 7.11317ptB^{n-1}v\notin{\rm Im}\,B^{n}.\end{array}\end{cases}$
@@
$\text{rk}\,(C^{n})=\begin{cases}\begin{array}[]{cc}\text{rk}\,(B^{n}),&\text{if}\hskip 7.11317ptB^{n-1}v\in{\rm Im}\,B^{n},\\ \text{rk}\,(B^{n})+1,&\text{if}\hskip 7.11317ptB^{n-1}v\notin{\rm Im}\,B^{n}.\end{array}\end{cases}$
@@
Throughout this article, we use the following notation. The vector $\epsilon_{\ell}$ denotes the $\ell$th vector of the canonical basis of ${\mathbb{R}}^{d}$, for $\ell=1,\dots,d$.
The idea of the proof of the Proposition 3.4 is contained in [2, Theorem 3.4].
Proof of Proposition 3.4.
Set $t=m_{1}+\cdots+m_{r}$ and we rename the elements of the canonical basis of $\mathbb{R}^{m}$ as $\{\epsilon^{1}_{1},\dots,\epsilon^{1}_{m_{1}},\dots,\epsilon_{1}^{r},\dots,\epsilon_{m_{r}}^{r},\epsilon_{t+1},\dots,\epsilon_{t+s}\}$. We can write $v$ as
@@
where $v_{i}\in\mathbb{R}$, $w\in{\rm Im}\,B$ and $w_{0}\in\text{span}\{\epsilon_{t+1},\ldots,\epsilon_{t+s}\}\subseteq\operatorname{Ker}\,B$. Note that by Lemma 3.5 we may assume $w=0$. We deal with the following cases:
$\begin{cases}{\rm(a)}\;v\in{\rm Im}\,B,\\ {\rm(b)}\;v\notin{\rm Im}\,B\text{ and }v_{i}\neq 0\text{ for some }i\in\{1,\dots,r\},\\ {\rm(c)}\;v\notin{\rm Im}\,B\text{ and }v\in\text{span}\{\epsilon_{t+1},\ldots,\epsilon_{t+s}\}.\end{cases}$
(a) If $v\in{\rm Im}\,B$, then by Lemma 3.5, $C$ is conjugate to $0_{1}\oplus B$.
(b) If $v\notin{\rm Im}\,B$ and there exists $i$ such that $v_{i}\neq 0$, set $d:=\text{min}\{i:v_{i}\neq 0\}$. Let
@@
In order to prove that $C$ and $A$ are conjugate, we show that $\dim(\operatorname{Ker}\,C^{n})=\dim(\operatorname{Ker}\,A^{n})$, for $n\in{\mathbb{N}}$. Moreover, since
$\dim(\operatorname{Ker}\,\mathcal{J}^{n}_{m_{i}})=\begin{cases}m_{i}&\text{if }n\geq m_{i},\\ n&\text{if }n<m_{i}.\end{cases}$
and $C$ and $A$ are both nilpotent matrices such that $C^{m_{1}+1}=A^{m_{1}+1}=0$, it is enough to show that $\dim(\operatorname{Ker}\,C^{n})=\dim(\operatorname{Ker}\,A^{n})$, for $1\leq n\leq m_{1}$. Since $v\notin{\rm Im}B$, in particular $v\neq 0$, and we have